OpenStudy (anonymous):

16x^2-240-32=0? Can someone check my answer? What am I doing wrong? I got( -1,2)

OpenStudy (anonymous):

16(x^2-15-2)=0 (x+1)(x-2)

OpenStudy (inkyvoyd):

You factored the equation incorrectly.

OpenStudy (anonymous):

how so?

OpenStudy (inkyvoyd):

(x+1)(x-2)=x^2-2x-2.

OpenStudy (inkyvoyd):

you'll have to use the quadratic equation for this one I believe, as b^2-4ac is not a perfect square.

OpenStudy (inkyvoyd):

*formula I mea

OpenStudy (anonymous):

I did- see this is what I had 16x^2-240-32x=0 then 16(x^2-15-2)=0 then divid each side by 16 then this is where I had problems

OpenStudy (inkyvoyd):

what do you mean exactly? that was correct, but 16(x^2-15-2)=0 x^2-15-2=0 at this point you factored the equation incorrectly. You must use the quadratic equation for this one.

OpenStudy (anonymous):

Thanks for your help. The answer was actually 3, -5